y=x+1
2x+y=10
how do you solve this using substitution?

Answer by Tobiasz(54)

y=x+1
2x+y=10

all you do is add part A into Part B , meaning plug it in and solve like a regular equation.
2x + x + 1 = 10
3x + 1 = 10 <--- one one from one and ten
3x = 9 <--- divide 3 from both sides
x+=+3
2(3)+y=10
6 + y = 10 <-- minus 6 from both sides
y+=+4
plug it in to check.!
2(3) + 4 = 10
6 + 4 = 10
10 = 10
CORRECT~!!